7.) Find X so that for any 3 × 3 real matrix A you get AX = XA = A [Hint : what property is being exhibited by real number p so that for any real w we get wp = pw = w then interpret for matrices.] 1 8.) Consider K =  1 −1 1 −1  then we get K2 = 0 Does this hold for real numbers? Motivate.


1
Expert's answer
2021-10-04T19:39:56-0400

17.

X="\begin{pmatrix}
 1&0&0 \\
 0&1&0\\ 
0&0&1
\end{pmatrix}"


Its equivalent to identify property of multiplication in real numbers where:

wp=pw=w given p= 1


Given K=

"\begin{bmatrix}
 1&-1;&1&-1 
 
\end{bmatrix}"


K2 ="\begin{pmatrix}
 1 & -1 \\
 1& -1
\end{pmatrix}" "\begin{pmatrix}
 1 & -1\\
 1& -1
\end{pmatrix}"

="\begin{pmatrix}
 0 & 0 \\
 0& 0
\end{pmatrix}"



The above is not true for real numbers.


Only a square of "0" will give a zero


0"\u00d7"0=0.



18.

-If ab=0, either a=0 or b=0

-Products of two non-zero numbers is always non-zero

But products of two non-zero matrices can be zero matrix


Using K given above


K="\begin{bmatrix}
 1&-1&;&1&-1 
 
\end{bmatrix}"


That is K"\begin{pmatrix}
 1 & -1 \\
 1 & -1
\end{pmatrix}"

K2=(K)(K)


="\begin{pmatrix}
 1 & -1 \\
 1& -1
\end{pmatrix}" "\begin{pmatrix}
 1 & -1\\
 1 & -1
\end{pmatrix}"


"\begin{pmatrix}
 1×1+-1×1 & 1×-1+-1×-1\\
 1×1+-1×1 & 1×-1+-1×-1
\end{pmatrix}"



="\begin{pmatrix}
 0 & 0\\
 0& 0
\end{pmatrix}"


=0 (Null matrix)